OpenStudy (anonymous):

what are the foci of the hyperbola given by the equation (x+10)^2/25-(y+1)^2/9=1

OpenStudy (agent0smith):

\[\frac{ (x+10)^2 }{ 25 }- \frac{ (y+1)^2 }{ 9 }=1\] Standard form of a horizontal hyperbola is:\[\frac{ (x-h)^2 }{ a^2 }- \frac{ (y-k)^2 }{b^2 }=1\]with foci \[(h \pm c, k)\] where c is...\[c^2 = a^2 +b^2\]

OpenStudy (anonymous):

So would it be -10 plus or mince sqrt(34,-1

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

That... looks correct!

OpenStudy (anonymous):

thank you so much!

OpenStudy (agent0smith):

You're welcome :)